Using Atomic Data Structures for Parallel Simulation

Computation Structures Group Memo 344 May, 1992

Paul S. Barth

In Proceddings of the Scalable High Performace Computing Conference, Williamsburg, VA, April 27,1992.

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988.